What is the density of brass at 60 degree Celsius, if the density at zero degree Celsius is 15.2 g/cm³ given linear expansivity of brass is 1.9*10^-5
1
Expert's answer
2020-05-07T10:06:37-0400
"\rho=\frac{\rho_0}{1+3\alpha\Delta T}"

"\rho=\frac{15.2}{1+3(1.9\cdot10^{-5})(60)}=15.15\frac{g}{cm^3}"